Complete solution of d^2y/dx^2 + 4y = cos 4x


1
Expert's answer
2021-05-07T14:37:07-0400

Let us complete solution of differential equation "\frac{d^2y}{dx^2} + 4y = \cos 4x."


Firstly, let us solve the characteristic equation of homogeneous differential equation:


"k^2+4=0"

"k_1=2i,\ k_2=-2i."


Therefore, the general solution of homogeneous equation is


"y=C_1\cos 2x + C_2\sin 2x."


Let us find the partial solution of non-homogeneous equation:


"y_p=A\cos 4x+B\sin 4x"


"y_p'=-4A\sin 4x+4B\cos 4x"


"y_p''=-16A\cos 4x-16B\sin 4x"


"-16A\cos 4x-16B\sin 4x+4A\cos 4x +4B\sin 4x=\cos 4x"


"-12A\cos 4x-12B\sin 4x=\cos 4x"


"-12A=1, -12B=0"


"A=-\frac{1}{12}, \ B=0."


Consequently the general solution of the differential equation "\frac{d^2y}{dx^2} + 4y = \cos 4x" is


"y=C_1\cos 2x + C_2\sin 2x-\frac{1}{12}\cos 4x."